# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v2

@@ 5.3 Impact of question length
The synthetic data questions in this study tend to be longer than the original human-written questions. The generating model is prompted to include all information required to disambiguate what the synthetic question is asking about, causing the questions to be longer and more detailed than the original dataset. Figure 5(a) highlights this trend. Some generating LMs take this to an extreme, causing the synthetic questions to become easier for LMs to answer correctly. We hypothesize this happens because the additional detail in the question places the LM under evaluation in the right “state of mind” so to speak, in the representation vector space. This effect is seen in Figure 5(b) where benchmark evaluation accuracy over-performance () is plotted. There is a rough trend where longer questions over-perform the expected trend line. It is worth noting that this effect does not show up in answer length (see Appendix B for detail and examples).
@@ 5.4 Benchmark question diversity
During synthetic question generation, the explicit topic extraction step provides a backstop against degenerate questions which only focus on the most obvious elements of a context chunk. To compare the variety and coverage of the generated questions we compute the average cosine similarity between all question pairs (excluding self pair) over the generated synthetic dataset. While most generating models produce similar diversity scores closely aligned with the human questions, some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 in Appendix C shows the question diversity per generating model along with the original question diversity.
@@ 5.5 Benchmark model accuracy correlation
Ideally, we want synthetic benchmarks to correlate well with the evaluation accuracy measured by human curated questions. This ends up being true for some generating models, but not all. Figure 6 shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the five bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 in Appendix E showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near evaluation accuracy (Pearson accuracy correlation ).
@@ 5.6 Benchmark model rank correlation
While direct Pearson correlation between human and synthetic benchmark evaluation accuracy is desirable, evaluations are often used to compare models. Therefore, a more important metric is the model rank order stability measured by Spearman rank correlation. Figure 2 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. Pearson accuracy correlation translates into Spearman ranking correlation and relative model order stability. Similarly, the overly easy gpt-4.1-mini synthetic benchmark in Figure 12 produces a mis-matched rank ordering (Spearman correlation ) in Figure 13(b). While ensembling various question generating LMs together will partially mitigate this effect, it is worth validating the set of generating models against the datasets in this paper to confirm the ensemble is not negatively impacting overall synthetic benchmark quality. The tendency of certain generating models to produce overly long, detailed questions that LMs find easier to answer may be the cause of the observed benchmark over-performance.
@@ 5.7 Multiple-choice vs open-ended questions
During this study, both multiple choice and open-ended questions were evaluated. None of our findings change whether using multiple choice questions for simplified evaluation benchmark grading, or open-ended questions to produce harder evaluation benchmarks where knowledge synthesis is required. The models that are good at creating synthetic multiple choice questions are also good at creating open ended questions. The models which create overly easy questions by including too much detail (like phi-4) do so for both open ended and MCQ, with the long question problem becoming even more pronounced for open ended questions as shown in Figure 9 in Appendix D.
Table 5 presents Pearson benchmark evaluation accuracy correlation and Spearman rank order correlation between the human and synthetic evaluations. See Appendix E for an expanded table showing correlations across all three LMs used for human question reformatting. Llama-4-Maverick was unexpectedly dominant in these final Table 5 results, even outperforming the ensemble of models (though not significantly).
@@
| Ensemble | 0.7452 | 0.7095 | 0.9679 | 0.9209 |
| Llama-3.2-3B-Instruct | 0.4857 | 0.4709 | 0.6143 | 0.7143 |
| Llama-3.3-70B-Instruct | 0.6615 | 0.6034 | 0.9571 | 0.8813 |
| Llama-3.1-8B-Instruct | 0.6608 | 0.6217 | 0.7964 | 0.7275 |
| Llama-4-Maverick | 0.7891 | 0.7233 | 0.9679 | 0.9648 |
| gemma-3-12b-it | 0.6879 | 0.4085 | 0.8929 | 0.8198 |
| gemma-3-27b-it | 0.7477 | 0.6962 | 0.9571 | 0.9516 |
| gemma-3-4b-it | 0.6575 | 0.5787 | 0.6393 | 0.9209 |
| gpt-4.1 | 0.6188 | 0.5216 | 0.9179 | 0.8549 |
| gpt-4.1-mini | 0.4015 | 0.5780 | 0.5643 | 0.8418 |
| o4-mini | 0.6542 | 0.5700 | 0.9250 | 0.7934 |
| phi-4 | 0.1547 | 0.4965 | 0.6964 | 0.1780 |
| Qwen2.5-7B-Instruct | 0.5813 | 0.5905 | 0.7321 | 0.8901 |
5.8 Case study on two academic papers
@@
To demonstrate our approach in a more realistic setting, we selected two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” [44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” [45]. The documents were chunked into 27 and 28 contexts respectively. “Current Solutions” averaged topics per context for a total of 210 questions. “Recent Advances” averaged topics per context for a total of 194 questions. For MCQ, the two synthetic benchmarks’ Spearman rank correlation is , for OE the rank correlation is . Under MCQ the gemma-3 models perform middle of the pack on these two benchmarks, but under OE questions the 12b and 27b outperform all other models. See Figure 14 in Appendix G for full rank ordering and examples of the generated questions.
@@ 5.9 Limitations
This work relies on a simple and possibly non-optimal document chunking approach. Extensions which cluster related topics or provide more fine-grained control over the information present in each chunk could improve the granularity and specificity of the generated benchmark. Including information found in tables, charts, and images would provide richer information sources that benchmark creation may take advantage of. Additionally, improving the nuance and difficulty of the questions significantly improves LM capability measurement. This likely requires careful management of the context information provided to the generating LM. Finally, the datasets used in this study to validate correlation with human written questions are all NLP extractive question answering type datasets. This limits the quality of the synthetic benchmarks as the grounding documents are fairly simplistic, especially compared to the two academic papers used in Section 5.8. Only the sec_qa dataset, which is grounded in an open source computer security textbook, comes close to the topic diversity and complexity of the example academic papers. All other validation datasets rely upon more general trivia type knowledge.
@@ 6 Conclusion
Our method supports automatic generation of both multiple-choice questions (optimizing for ease of grading) and open-ended questions (optimizing for diagnostic relevance), allowing users to choose based on their evaluation goals. Validation against six human-curated NLP dataset achieves a Spearman rank correlation of and a benchmark Pearson accuracy correlation of between synthetic and human curated questions, suggesting that our approach is consistent with human-generated question/answer benchmarks. Additionally, we identify a potential synthetic data generation failure mode: very long questions can artificially inflate model performance, likely by overspecifying information related to the answer. This methodology enables the creation of targeted, relevant benchmark evaluations of domain-specific Language Model capabilities, facilitating deeper exploration into the capabilities and limitations of Language Models deployed for complex knowledge work.
@@ Appendix A Expanded related works
This work draws on the rich history of language model evaluation, where LMs are already being used to judge and score the results of human curated benchmarks. When scoring outputs of an LM producing a multiple choice answer, it is often useful to have another LM reason about and extract the specified [A,B,C,D] value, instead of hoping the model under evaluation only produced the exact match answer letter [3, 4, 46]. This process becomes even more important when evaluating short answers from LMs instead of multiple choice responses, or when scoring the responses and providing reasoning as to why an answer was correct or incorrect [5, 6, 7]. This idea of automating question answer generation has been approached before. The most similar contemporary work is YourBench by Shashidhar et al. [8], which leverages the convergent idea of grounding documents supporting LM benchmark generation, demonstrating the synthetic replication of MMLU. Additionally, AutoBencher by Li et al. explores leveraging synthetic data (supported by privileged information) for model evaluation with RL to optimize the generated evaluation for discovering gaps in LM capability [9]. The use of grounding documents or privileged information greatly improves synthetic data benchmark utility, as the evaluation methodology no longer relies on simply asking an LM for questions (and answers).
@@
Early Question Answer Generation (QAG) approaches attempted to solve the question creation from a context + correct answer input. These approaches include improvements in selecting better LM samples [47], increasing sample diversity over distributional shifts [48], explorations of interrogative prompting against the context [49], and evaluating round trip information consistency [50]. These early methods sometimes fine-tune generation models to improve task performance, though largely that is only done with smaller models [51, 52]. Some work focused on fine-tuning specialty models to improve option generation for multiple-choice questions [52]. Other early work during the shift to general purpose LMs involved leveraging those language models directly as data generation sources. This took the form of prompting the model to write questions (without grounding context) [53, 54, 55, 56], or configuring an LM-as-evaluator/interrogator dynamically interacting with the LM under evaluation to explore capability in a reference-free manner [53]. These approaches attempt to address the dataset leakage problem where benchmarks, once published, become part of the training dataset for the next round of models. As the capability of LMs increases over time, setting up adversarial LM-evaluator vs LM-evaluatee systems is becoming more feasible [57]. Nevertheless, extending those systems such that the LM-evaluator has grounding documents to pull from should strengthen the performance.
@@
With the rise of general purpose token-tumblers like ChatGPT, longer and more complex grounding documents become increasingly useful. There is no longer a need to constrain synthetic benchmark creation to a short paragraph context + correct answer input; the LM can operate on larger and more complex document chunks, with less pre-processing required to make the input context data useful. Contemporaneous work YourBench [8] and AutoBencher [9] can leverage lightly-filtered web-pages as the context, chunking the data into rough paragraphs for ingestion into the question generation and grounding system. The SecQA benchmark was created partly through synthetic data grounded in an open-source computer security textbook [10]. Similarly (though much earlier), the PubMedQA benchmark was created through human and synthetic data generation grounded in biomedical research paper abstracts [11]. Some research demonstrates that operating purely on document summaries produces inferior generated questions compared to leveraging the full document [12]. With modern long-context LMs, summarization is becoming less of a problem, as models can manipulate significant volumes of text within their attention context. Synthetic data question generation systems have seen rapid improvement. This is partly driven by capability improvements in the LMs powering synthetic data generation. However, there has also been exploration into improvements in prompting for diverse question generation [13] and breadth of topic generation for chatbot arena style questions [14]. Enhancing retrieval augmented generation (RAG) systems for domain specific applications has also used synthetic data question generation to compare baseline RAG to fine-tuning [15].
@@
There are many aspects of text that one might want to quantify, however imprecisely. However, devoting the required copious human annotation hours is just impractical. Current LMs end up being acceptable feature extractors, when provided detailed prompts. Chiang and Lee demonstrate that LMs can substitute for human evaluators in assessing text quality with strong correlation in the results [16]. Similarly, Han et al. leverage LM evaluators for judging quality and find that results correlate highly with human judgments [17]. The use of LMs to extract features from text is not to dissimilar from using the LM-as-a-Judge [18] pattern for large scale open ended capability evaluations [18, 19, 21]. The LM is capable of grading open ended evaluation benchmark questions for correctness; therefore determining features other than answer correctness from a text sample is viable. However, there are limits on what LM-as-a-Judge is capable of ground truthing, especially in fields like information retrieval [20, 21] where LMs under-perform on more complex relevance judgments. Relying on LMs to extract properties like answer correctness given sufficient scaffolding and context is a well-used pattern.

### v3

@@ 5.3 Impact of question length
The synthetic data questions in this study tend to be longer than the original human-written questions. The generating model is prompted to include all information required to disambiguate what the synthetic question is asking about, causing the questions to be longer and more detailed than the original dataset. Figure˜6(a) highlights this trend. Some generating LMs take this to an extreme, causing the synthetic questions to become easier for LMs to answer correctly. We hypothesize this happens because the additional detail in the question places the LM under evaluation in the right “state of mind” so to speak, in the representation vector space. This effect is seen in Figure˜6(b) where benchmark evaluation accuracy over-performance () is plotted. There is a rough trend where longer questions over-perform the expected trend line. It is worth noting that this effect does not show up in answer length (see Appendix˜C for detail and examples).
@@ 5.4 Benchmark question diversity
During synthetic question generation, the explicit topic extraction step provides a backstop against degenerate questions which only focus on the most obvious elements of a context chunk. To compare the variety and coverage of the generated questions we compute the average cosine similarity between all question pairs (excluding self pair) over the generated synthetic dataset. While most generating models produce similar diversity scores closely aligned with the human questions, some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure˜10 in Appendix˜D shows the question diversity per generating model along with the original question diversity.
@@ 5.5 Benchmark model accuracy correlation
Ideally, we want synthetic benchmarks to correlate well with the evaluation accuracy measured by human curated questions. This ends up being true for some generating models, but not all. Figure˜7 shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the five bolded ensemble models in Table˜2. This ensemble produces a Pearson accuracy correlation of between the evaluation accuracies for synthetic and human benchmarks (Figure˜15(b)). Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure˜14 in Appendix˜F showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near evaluation accuracy (Pearson accuracy correlation ).
@@ 5.6 Benchmark model rank correlation
While direct Pearson correlation between human and synthetic benchmark evaluation accuracy is desirable, evaluations are often used to compare models. Therefore, a more important metric is the model rank order stability measured by Spearman rank correlation. Figure˜3 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. Pearson accuracy correlation translates into Spearman ranking correlation and relative model order stability. Similarly, the overly easy gpt-4.1-mini synthetic benchmark in Figure˜14 produces an odd rank ordering (Spearman correlation ) in Figure˜15(b) with a very flat accuracy curve. While ensembling various question generating LMs together will partially mitigate this effect, it is worth validating the set of generating models against the datasets in this paper to confirm the ensemble is not negatively impacting overall synthetic benchmark quality. The tendency of certain generating models to produce overly long, detailed questions that LMs find easier to answer may be the cause of the observed benchmark over-performance.
@@ 5.7 Multiple-choice vs open-ended questions
During this study, both multiple choice and open-ended questions were evaluated. None of our findings change whether using multiple choice questions for simplified evaluation benchmark grading, or open-ended questions to produce harder evaluation benchmarks where knowledge synthesis is required. The models that are good at creating synthetic multiple choice questions are also good at creating open ended questions. The models which create overly easy questions by including too much detail (like phi-4) do so for both open ended and MCQ, with the long question problem becoming even more pronounced for open ended questions as shown in Figure˜11 in Appendix˜E.
Table˜5 presents Pearson benchmark evaluation accuracy correlation and Spearman rank order correlation between the human and synthetic evaluations. Llama-4-Maverick was unexpectedly dominant for multiple-choice in these final Table˜5 results, even outperforming the ensemble of models. Similarly gemma-3-23b-it over-performed the ensemble for open-ended questions.
@@
| Ensemble | 0.7434 | 0.7238 | 0.9066 | 0.6099 |
| Llama-3.2-3B-Instruct | 0.5444 | 0.7288 | 0.2527 | 0.4176 |
| Llama-3.3-70B-Instruct | 0.6807 | 0.6992 | 0.8626 | 0.4725 |
| Llama-3.1-8B-Instruct | 0.5484 | 0.7402 | 0.8901 | 0.4780 |
| Llama-4-Maverick | 0.8232 | 0.7833 | 0.9231 | 0.6758 |
| gemma-3-12b-it | 0.6879 | 0.6521 | 0.7363 | 0.5000 |
| gemma-3-27b-it | 0.7325 | 0.8192 | 0.8846 | 0.8516 |
| gemma-3-4b-it | 0.6495 | 0.7771 | 0.5549 | 0.7308 |
| gpt-4.1 | 0.5784 | 0.5071 | 0.7967 | 0.4341 |
| gpt-4.1-mini | 0.2759 | 0.2505 | 0.7527 | 0.4560 |
| o4-mini | 0.6687 | 0.5290 | 0.8187 | 0.5495 |
| phi-4 | 0.3489 | 0.3936 | 0.5989 | 0.4725 |
| Qwen2.5-7B-Instruct | 0.5901 | 0.6018 | 0.4286 | 0.6099 |
| Qwen3-235B-A22B-Instruct-2507 | 0.6084 | 0.6017 | 0.7857 | 0.5495 |
5.8 Case study on academic papers
@@
To demonstrate our approach in a more realistic setting, we selected three documents (two of which are post-LM training knowledge cutoff): “Current Solutions and Future Trends for Robotic Prosthetic Hands” (Mendez et al., 2021), “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (Chen et al., 2025), and the recently released “America’s AI Action Plan” (4). The documents were chunked into roughly 30 chunks each. “Current Solutions” averaged topics per context for a total of 210 questions. “Recent Advances” averaged topics per context for a total of 194 questions. “AI Action Plan” produced a total of 235 questions. For MCQ, the synthetic benchmarks’ Spearman rank correlation is , for OE the rank correlation is . These correlations reflect the trend of overall model capability, as both source documents are from post training cutoff, so models forced to puzzle out what the likely answers are. Unlike earlier results where the correlation was a comparison to the human curated questions, these correlations only compare the LM benchmark performance between the two academic papers. Under MCQ the gemma-3 models perform middle of the pack on these two benchmarks, but under OE questions the 12b and 27b outperform all other models. See Figure˜16 in Appendix˜H for full rank ordering and examples of the generated questions.
5.9 Replicating a subset of MMLU-Pro
The datasets used herein to validate the generative benchmarking approach were never designed to evaluate general LM knowledge. To alleviate this weakness and demonstrate the capability of this generative benchmark creation approach, we replicated a subset of the MMLU-Pro benchmark (Wang et al., 2024b) using creative commons textbooks from OpenStax (Stafford and Flatley, 2018). Six of the categories from MMLU-Pro were replicated by using a related textbook as the grounding context: biology (Clark et al., 2018), chemistry (Flowers et al., 2019), computer science (Franchitti, 2024), math (Abramson, 2021), physics (Urone and Hinrichs, 2022), and psychology (Spielman et al., 2020). We used gpt-oss-120b to construct the generative benchmarks. Each textbook was converted into markdown, chunked into 4k token blocks, and ingested into the generative benchmarking pipeline. The resulting generative benchmarks were evaluated using the full suite of models in Table˜2. Figure˜8 demonstrates the correlation between the generative replication of each MMLU-pro subset and the original human-curated benchmark results.
The generative replication of MMLU-pro benchmark produces a Pearson accuracy correlation of 0.78 with the original benchmark (Spearman rank correlation of 0.83). Given that the generated questions were created using only relevant textbooks where question topics do not align, the resulting level of agreement between human curated and grounded generative benchmarking is surprising. Additionally, the accuracy aligns well with underlying model capability as measured by the human-curated MMLU-Pro benchmark.
To validate the quality of the generative questions, the authors manually reviewed 120 questions (20 each randomly selected from each MMLU-pro subset). This evaluation probed two elements, the question/answer pair correctness and grounding in the source document context. We found that 111 of 120 questions had completely accurate generative question/answer pairs, meaning the question was well formed and the answer was correct. This yields a noise floor for generative benchmarking approach of approximately . 104 of 120 questions had fully supported grounding in the document. A strong majority of the grounding failures fell into two categories: 1) the question writing model adding relevant correct information that was missing from the document, or 2) the generative question referencing the document and not being fully independent.
An example high quality generative question from the physics subset is as shown:
We also provide an example of a question where the system failed to produce a correct answer:
A more detailed analysis of the human-evaluated question set is provided in Appendix˜B. The MMLU-pro replication used gpt-oss-120b for question generation. It is possible that using a stronger frontier model would reduce the generation and grounding error rate observed.
@@ 5.10 Limitations
This work relies on a simple and possibly non-optimal document chunking approach. Extensions which cluster related topics or provide more fine-grained control over the information present in each chunk could improve the granularity and specificity of the generated benchmark. Including information found in tables, charts, and images would provide richer information sources that benchmark creation may take advantage of. Additionally, improving the nuance and difficulty of the questions significantly improves LM capability measurement. This likely requires careful management of the context information provided to the generating LM. This approach is not yet suitable for evaluating the absolute frontier of model capability, but many applications rely on weaker models with lower inference costs. Finally, except for the subset of MMLU-Pro replicated using Creative Commons textbooks, the datasets used in this study to validate correlation with human written questions are all NLP extractive question answering type datasets. This limits the quality of the synthetic benchmarks as the grounding documents are fairly simplistic, especially compared to the two academic papers used in Section˜5.8. Only the PubMedQA and sec_qa datasets come close to the topic diversity and complexity of the example academic papers. All other validation datasets rely upon more general factual knowledge.
@@ 6 Conclusion
The majority of the invalid question/answer pairs stemming from either missing information or grounding failures could be rectified by an adversarial verifier running within the question generation loop. In future work we plan to explore this research direction both for generative benchmarking and dynamic context grounding.
Our method supports automatic generation of both multiple-choice questions (optimizing for ease of grading) and open-ended questions (optimizing for diagnostic relevance), allowing users to choose based on their evaluation goals. Validation against human-curated NLP dataset achieves an ensemble Spearman rank correlation of and a benchmark Pearson accuracy correlation of between synthetic and human curated questions, suggesting that our approach is consistent with human-generated question/answer benchmarks. Additionally, we identify a potential synthetic data generation failure mode: very long questions can artificially inflate model performance, likely by overspecifying information related to the answer. This methodology enables the creation of targeted, relevant benchmark evaluations facilitating deeper exploration into the domain specific capabilities and limitations of Language Models deployed for complex knowledge work.
@@ Appendix A Expanded related works
This work draws on the rich history of language model evaluation, where LMs are already being used to judge and score the results of human curated benchmarks. When scoring outputs of an LM producing a multiple choice answer, it is often useful to have another LM reason about and extract the specified [A,B,C,D] value, instead of hoping the model under evaluation only produced the exact match answer letter (Samuylova, 2024; Roucher, 2024; Wolfe, 2024). This process becomes even more important when evaluating short answers from LMs instead of multiple choice responses, or when scoring the responses and providing reasoning as to why an answer was correct or incorrect (Ip, 2024; Oh et al., 2024; Cook et al., 2024). This idea of automating question answer generation has been approached before. The most similar contemporary work is YourBench by Shashidhar et al. (Shashidhar et al., 2025), which leverages the convergent idea of grounding documents supporting LM benchmark generation, demonstrating the synthetic replication of MMLU. Additionally, AutoBencher by Li et al. explores leveraging synthetic data (supported by privileged information) for model evaluation with RL to optimize the generated evaluation for discovering gaps in LM capability (Li et al., 2025b). The use of grounding documents or privileged information greatly improves synthetic data benchmark utility, as the evaluation methodology no longer relies on simply asking an LM for questions (and answers).
@@
Early Question Answer Generation (QAG) approaches attempted to solve the question creation from a context + correct answer input. These approaches include improvements in selecting better LM samples (Yuan et al., 2022), increasing sample diversity over distributional shifts (Shinoda et al., 2020), explorations of interrogative prompting against the context (Eo et al., 2023), and evaluating round trip information consistency (Alberti et al., 2019). These early methods sometimes fine-tune generation models to improve task performance, though largely that is only done with smaller models (Ushio et al., 2023; Zhou, 2024). Some work focused on fine-tuning specialty models to improve option generation for multiple-choice questions (Zhou, 2024). Other early work during the shift to general purpose LMs involved leveraging those language models directly as data generation sources. This took the form of prompting the model to write questions (without grounding context) (Bai et al., 2023; Tan et al., 2024; Perez et al., 2023; Zhang and Wang, 2022), or configuring an LM-as-evaluator/interrogator dynamically interacting with the LM under evaluation to explore capability in a reference-free manner (Bai et al., 2023). These approaches attempt to address the dataset leakage problem where benchmarks, once published, become part of the training dataset for the next round of models. As the capability of LMs increases over time, setting up adversarial LM-evaluator vs LM-evaluatee systems is becoming more feasible (Li et al., 2025a). Nevertheless, extending those systems such that the LM-evaluator has grounding documents to pull from should strengthen the performance.
@@
With the rise of general purpose token-tumblers like ChatGPT, longer and more complex grounding documents become increasingly useful. There is no longer a need to constrain synthetic benchmark creation to a short paragraph context + correct answer input; the LM can operate on larger and more complex document chunks, with less pre-processing required to make the input context data useful. Contemporaneous work YourBench (Shashidhar et al., 2025) and AutoBencher (Li et al., 2025b) can leverage lightly-filtered web-pages as the context, chunking the data into rough paragraphs for ingestion into the question generation and grounding system. The SecQA benchmark was created partly through synthetic data grounded in an open-source computer security textbook (Liu, 2023). Similarly (though much earlier), the PubMedQA benchmark was created through human and synthetic data generation grounded in biomedical research paper abstracts (Jin et al., 2019). Some research demonstrates that operating purely on document summaries produces inferior generated questions compared to leveraging the full document (Bhat et al., 2023). With modern long-context LMs, summarization is becoming less of a problem, as models can manipulate significant volumes of text within their attention context. Synthetic data question generation systems have seen rapid improvement. This is partly driven by capability improvements in the LMs powering synthetic data generation. However, there has also been exploration into improvements in prompting for diverse question generation (Yadav et al., 2024) and breadth of topic generation for chatbot arena style questions (Li et al., 2024). Enhancing retrieval augmented generation (RAG) systems for domain specific applications has also used synthetic data question generation to compare baseline RAG to fine-tuning (Balaguer et al., 2024).
@@
There are many aspects of text that one might want to quantify, however imprecisely. However, devoting the required copious human annotation hours is just impractical. Current LMs end up being acceptable feature extractors, when provided detailed prompts. Chiang and Lee demonstrate that LMs can substitute for human evaluators in assessing text quality with strong correlation in the results (Chiang and Lee, 2023). Similarly, Han et al. leverage LM evaluators for judging quality and find that results correlate highly with human judgments (Han et al., 2024). The use of LMs to extract features from text is not to dissimilar from using the LM-as-a-Judge (Xu et al., 2023) pattern for large scale open ended capability evaluations (Xu et al., 2023; Zheng et al., 2023; Szymanski et al., 2025). The LM is capable of grading open ended evaluation benchmark questions for correctness; therefore determining features other than answer correctness from a text sample is viable. However, there are limits on what LM-as-a-Judge is capable of ground truthing, especially in fields like information retrieval (Soboroff, 2025; Szymanski et al., 2025) where LMs under-perform on more complex relevance judgments. Relying on LMs to extract properties like answer correctness given sufficient scaffolding and context is a well-used pattern.